# SGDFuse: SAM-Guided Diffusion for High-Fidelity Infrared and Visible Image Fusion

### v2

@@ Abstract
Infrared and visible image fusion (IVIF) aims to combine the thermal radiation information from infrared images with the rich texture details from visible images to enhance perceptual capabilities for downstream visual tasks. However, existing methods often fail to preserve key targets due to a lack of deep semantic understanding of the scene, while the fusion process itself can also introduce artifacts and detail loss, severely compromising both image quality and task performance. To address these issues, this paper proposes SGDFuse, a conditional diffusion model guided by the Segment Anything Model (SAM), to achieve high-fidelity and semantically-aware image fusion. The core of our method is to utilize high-quality semantic masks generated by SAM as explicit priors to guide the optimization of the fusion process via a conditional diffusion model. Specifically, the framework operates in a two-stage process: it first performs a preliminary fusion of multi-modal features, and then utilizes the semantic masks from SAM jointly with the preliminary fused image as a condition to drive the diffusion model’s coarse-to-fine denoising generation. This ensures the fusion process not only has explicit semantic directionality but also guarantees the high fidelity of the final result. Extensive experiments demonstrate that SGDFuse achieves state-of-the-art performance in both subjective and objective evaluations, as well as in its adaptability to downstream tasks, providing a powerful solution to the core challenges in image fusion. The code of SGDFuse is available at https://github.com/boshizhang123/SGDFuse.
@@ 1 Introduction
Despite advancements in feature extraction and fusion strategies, inherent design limitations in existing methods lead to a core problem: a ’semantic blindness’. For instance, methods based on CNNs are constrained by their local receptive fields, an architectural limitation that hinders their ability to comprehend global scene structures, often resulting in blurred boundaries for large-scale targets or inconsistent background luminance. On the other hand, while GAN-based approaches can produce visually sharp images, their training relies on pixel-level fidelity constraints (e.g., intensity and gradient losses), which struggle to capture complex cross-modal semantic relationships. This not only causes training instability but also frequently leads to the erroneous suppression of salient thermal targets in infrared images. These specific design flaws point to a fundamental issue: the models are unable to semantically differentiate between foreground targets and background textures, thus failing to perform selective information enhancement. This ”can’t see the forest for the trees” paradigm directly causes a cascade of problems, including blurred target boundaries, loss of critical structures, and erroneous suppression of thermal features. Ultimately, this semantic deficiency is the fundamental bottleneck constraining the efficacy of fused images in downstream tasks like object detection and semantic segmentation. As illustrated in Fig. 1, the failure of existing models to preserve task-critical target regions severely undermines the practical utility of fusion technology.
Recent advancements in large-scale vision models offer new promise [12]. The Segment Anything Model (SAM) [13] represents a key breakthrough owing to its outstanding semantic segmentation capabilities. The high-quality semantic masks it generates can provide multi-modal images with rich semantic priors (as shown in Fig. 2). This wealth of semantic information can significantly alleviate the ”semantic blindness” inherent in existing methods. It guides the model to more effectively preserve the thermal radiation features and visible details of critical targets, thereby achieving a more semantically consistent fused representation. On the other hand, image fusion, as an image generation task, is typically accompanied by processes of information compression and modal reconstruction [14]. Although this process can enhance visual consistency, it often leads to a degradation in detail quality, manifesting as blurred textures, indistinct boundaries, and visual artifacts. Such degradation not only diminishes the perceptual quality of the fused image but also significantly impairs the performance of downstream tasks that rely on structural information, such as object detection and semantic segmentation. Diffusion models, by virtue of their
@@
iterative denoising generation mechanism, can stably produce high-fidelity, artifact-free images, making them exceptionally well-suited for the demands of IVIF, namely detail restoration and modal balance. More importantly, when integrated with high-quality semantic priors, diffusion models can perform semantically-aware conditional generation. This enables the fusion process to be more targeted in its representation of structures, textures, and target regions, significantly enhancing its adaptability for downstream tasks.
@@
Building upon the aforementioned insights, we propose the SAM-Guided Diffusion Fusion Network (SGDFuse), which employs a two-stage training framework to achieve synergistic optimization of structural semantic understanding and high-fidelity image generation. The first stage focuses on multi-modal feature extraction and fusion. We design a Multi-Scale Feature Enhancement Module (MSFEM) and a Transformer Block to model the local structures of the infrared image and the global semantics of the visible image, respectively. A preliminary fused image is then generated through the effective alignment of these cross-modal features. The second stage is oriented towards task-oriented optimization and high-fidelity reconstruction. Here, we construct a fused image optimization mechanism based on a conditional diffusion model, concentrating on structural detail restoration and task-relevant semantic enhancement. Specifically, we leverage SAM to generate rich semantic prior masks for both the infrared and visible images. These masks are concatenated with the preliminary fused image to form a five-channel, task-aware input. This input guides the conditional diffusion process to progressively reconstruct the image structure and enhance semantic details. Concurrently, we design a Hierarchical Feature Aggregation Head (HFAH) to aggregate deep semantic information from multiple scales, further improving the fused image’s representation of target structures and regional boundaries. Through the organic integration of semantic guidance and high-fidelity reconstruction, our method establishes an effective synergy between structural understanding and detail generation. This significantly boosts the semantic integrity, visual quality, and adaptability for downstream tasks of the resulting fused image.
@@
    Semantic-Aware Fusion: By utilizing high-quality semantic masks from SAM as explicit priors, our method effectively overcomes the ”semantic blindness” of conventional approaches. This results in a superior preservation and enhancement of key semantic information, like infrared thermal targets and visible textures, boosting the semantic integrity of the fused image.
@@
    High-Fidelity Image Optimization Mechanism: We pioneer the use of a conditional diffusion model to perform high-fidelity optimization on the fused image. By virtue of their robust training stability and excellent generative power, diffusion models enable a high-precision reconstruction of the preliminary fused information, effectively minimizing artifacts and ensuring the final representation maintains maximum fidelity under semantic guidance.
@@
    Two-Stage Task-Oriented Fusion Framework: We propose a framework that innovatively unifies structural understanding and high-quality generation by synergistically integrating multi-modal feature fusion with task-aware, diffusion-based optimization. This approach, guided by semantic priors, effectively boosts the fused image’s performance on downstream tasks and its representation of regional structures.
@@ 2.2 SAM and Its Applications
The Segment Anything Model (SAM) [13], as a powerful pretrained model, has become a current research hotspot in the field of image processing due to its zero-shot generalization capability in image segmentation and object detection tasks [24]. SAM can provide high-precision semantic masks across various vision tasks and has been widely applied in low-level vision tasks such as dehazing [25], and low-light image enhancement [26], as well as high-level tasks like object detection [27], demonstrating outstanding performance. By deeply understanding image content, SAM effectively captures fine details and accurately delineates target regions. Therefore, SAM is particularly suitable for infrared and visible image fusion tasks, we leverage SAM-generated segmentation masks as explicit, high-level semantic priors to guide the fusion process.
@@ 3.1 Motivation
Our primary and core motivation is to inject explicit semantic understanding capabilities into the fusion process. We are no longer satisfied with the indiscriminate processing of low-level features by existing methods, but rather seek a novel mechanism that enables the model to truly ”understand” the image and distinguish ”target” from ”background”. For this, we introduce SAM. The high-quality semantic masks it generates can provide the fusion process with unprecedented, strong semantic priors concerning scene structure and object boundaries, which becomes our ideal tool for solving the core problem of ”semantic blindness”. However, high-level semantic guidance alone is insufficient. To accurately translate this precise semantic intent—for instance, ’to preserve the thermal features of a target while enhancing the texture of the background’—into a high-quality, pixel-level output, one must rely on an equally powerful high-fidelity generative framework. Traditional reconstruction methods are prone to generating artifacts and detail loss when handling such complex conditions, which severely undermines the efficacy of semantic guidance. Therefore, our second motivation is to employ a Diffusion Model, renowned for its superior generative capabilities, as the backbone for image reconstruction. This ensures that the high-level semantic commands
@@
can be rendered losslessly and with maximum fidelity.
The core of our paradigm is the Deep Coupling of semantic understanding and image generation, rather than treating them as isolated steps. We leverage high-quality semantic masks from SAM as strong conditional priors to guide the iterative denoising of a conditional diffusion model. This endows the generation process with Precise Semantic Directivity, enabling the model to intelligently preserve thermal targets while reconstructing texture details based on semantic importance. Ultimately, SGDFuse unifies the contradictory objectives of semantic integrity and visual quality, producing fused images that are both visually compelling and highly effective for downstream tasks.
3.2 Network Architecture
In this section, we provide a detailed description of the proposed image fusion framework, SGDFuse. This fusion framework adopts a two-stage architecture, as illustrated in Fig. 3, aiming to generate high-quality fused images with structural fidelity and semantic consistency.
@@ 3.3 Multi-Scale Feature Enhancement Module (MSFEM)
As shown in Fig. 3(c), the MSFEM employs a multi-branch parallel structure, where each branch utilizes convolutional operations with different receptive fields (e.g., $\times1$, $\times3$, $\times5$, $\times7$) to capture structural features ranging from local to medium scales. Specifically, given an input image, feature representations are first extracted separately by four sets of convolutional kernels:
@@ 3.5 Hierarchical Feature Aggregation Head (HFAH)
To further enhance feature representation in the decoding stage, we introduce a Hierarchical Feature Aggregation Head (HFAH) into the decoder path of the denoising network. HFAH integrates multi-level decoded features and incorporates a spatial attention mechanism to jointly optimize structural detail and semantic consistency. Specifically, all weighted hierarchical features are concatenated and passed through a fusion head to generate the final fused image $I_{f}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$, which exhibits clearer structural boundaries and improved detail preservation. The fusion head consists of multiple $\times3$ convolutional layers designed to map the high-dimensional, multi-scale diffusion features into a final three-channel fused image. A Tanh activation function is applied at the output to enhance texture continuity and fine detail expression. Benefiting from HFAH’s precise modeling of structural edges and semantic regions, this module significantly improves the boundary fidelity and fine-grained reconstruction quality of the final fused image.
@@ 3.6.2 Stage II: Diffusion Optimization Training
4 EXPERIMENTS
4.1 Experimental Settings
4.1.1 Datasets
[…]
4.1.2 Evaluation Metrics
4.1.3 Training Details
4.2 Generalization Comparison Experiments
To evaluate the generalization capability of our network, we compare it against the following representative image fusion methods: U2Fusion [1], TarDAL [38], PIAFusion [10], IGNet [27], CDDFuse [32], Dif-fusion [5], PFCFuse [34], SHIP [24], EMMA [28], Text-DiFuse [11], DAFusion [15], MaeFuse [6], and SAGE [12].
4.2.1 Results and Analysis on the MSRS Dataset
Fig. 5 shows qualitative comparisons under diverse conditions, including varying illumination, multiple targets, and day/night scenes. While U2Fusion and PLAFusion exhibit blurred details and brightness inconsistency, IGNet, CDDFuse, and MaeFuse partially enhance thermal targets but often lack texture fidelity or introduce artifacts. In contrast, our method effectively preserves thermal saliency, enhances structural clarity, and maintains global luminance balance, achieving superior perceptual quality and semantic consistency.
4.2.2 Results and Analysis on the M3FD Dataset
Fig. 6 shows qualitative comparisons on challenging scenes from the M3FD dataset, including nighttime roads, uneven lighting, low-contrast IR targets, and complex textures. Methods like U2Fusion and TarDAL suffer from blurred structures and missing textures, while IGNet, PLAFusion, and CDDFuse offer partial improvements but still lack edge continuity and luminance balance. MaeFuse and SAGE improve global coherence but exhibit limitations in detail preservation. In contrast, our method consistently restores thermal contours and visible textures with natural brightness and clear contrast. Especially in scenes with weak IR signals and cluttered backgrounds, SGDFuse achieves accurate foreground–background separation, highlighting its superiority in structural refinement and semantic consistency.

### v4

@@ Abstract
Infrared and visible image fusion (IVIF) aims to combine the thermal radiation information from infrared images with the rich texture details from visible images to enhance perceptual capabilities for downstream visual tasks. However, existing methods often fail to preserve key targets due to a lack of deep semantic understanding of the scene, while the fusion process itself can also introduce artifacts and detail loss, severely compromising both image quality and task performance. To address this, this paper introduces SGDFuse, a novel semantic-guided generative framework. We pioneer a new paradigm that moves beyond traditional low-level feature mapping by deeply coupling high-level semantic priors from the Segment Anything Model (SAM) with the high-fidelity generative power of a conditional diffusion model. The core of our method is a synergistic two-stage process: it first performs a preliminary fusion to establish a robust structural foundation, and then utilizes semantic masks from SAM, jointly with this preliminary image, as a strong condition to steer the diffusion model’s coarse-to-fine refinement. This ensures the fusion process not only has explicit semantic directionality but also guarantees the high fidelity of the final result. Extensive experiments demonstrate that SGDFuse achieves state-of-the-art performance in both subjective and objective evaluations, as well as in its adaptability to downstream tasks, providing a powerful solution to the core challenges in image fusion. The code of SGDFuse is available at https://github.com/boshizhang123/SGDFuse.
@@ 1 Introduction
Despite advancements in feature extraction and fusion strategies, inherent design limitations in existing methods lead to a core problem: a ’semantic blindness’. For instance, methods based on CNNs are constrained by their local receptive fields, an architectural limitation that hinders their ability to comprehend global scene structures, often resulting in blurred boundaries for large-scale targets or inconsistent background luminance. On the other hand, while GAN-based approaches can produce visually sharp images, their training relies on pixel-level fidelity constraints (e.g., intensity and gradient losses), which struggle to capture complex cross-modal semantic relationships. This not only causes training instability but also frequently leads to the erroneous suppression of salient thermal targets in infrared images. These specific design flaws point to a fundamental issue: the models are unable to semantically differentiate between foreground targets and background textures, thus failing to perform selective information enhancement. This ”can’t see the forest for the trees” paradigm directly causes a cascade of problems, including blurred target boundaries, loss of critical structures, and erroneous suppression of thermal features. Ultimately, this semantic deficiency is the fundamental bottleneck constraining the efficacy of fused images in downstream tasks like object detection and semantic segmentation. As illustrated in Figure 1, the failure of existing models to preserve task-critical target regions severely undermines the practical utility of fusion technology.
Recent advancements in large-scale vision models offer new promise [12]. The Segment Anything Model (SAM) [13] represents a key breakthrough owing to its outstanding semantic segmentation capabilities. The high-quality semantic masks it generates can provide multi-modal images with rich semantic priors (as shown in Figure 2). This wealth of semantic information can significantly alleviate the ”semantic blindness” inherent in existing methods. It guides the model to more effectively preserve the thermal radiation features and visible details of critical targets, thereby achieving a more semantically consistent fused representation. On the other hand, image fusion, as an image generation task, is typically accompanied by processes of information compression and modal reconstruction [14]. Although this process can enhance visual consistency, it often leads to a degradation in detail quality, manifesting as blurred textures, indistinct boundaries, and visual artifacts. Such degradation not only diminishes the perceptual quality of the fused image but also significantly impairs the performance of downstream tasks that rely on structural information, such as object detection and semantic segmentation. Diffusion models, by virtue of their iterative denoising generation mechanism, can stably produce high-fidelity, artifact-free images, making them exceptionally well-suited for the demands of IVIF, namely detail restoration and modal balance. More importantly, when integrated with high-quality semantic priors, diffusion models can perform semantically-aware conditional generation. This enables the fusion process to be more targeted in its representation of structures, textures, and target regions, significantly enhancing its adaptability for downstream tasks.
@@
Building upon the aforementioned insights, we argue that the field of image fusion must undergo a conceptual shift-moving away from traditional ’pixel reorganization’ or ’feature mapping’ tasks towards what we propose as a new ’Semantic-Guided Generation’ (SGG) paradigm. The core insight of the SGG paradigm is that a high-fidelity and semantically-faithful fused image should not be reorganized from source pixels. Instead, it must be generated anew by a powerful generative model, and this generation process must be explicitly controlled and steered by high-level semantic priors. In this paper, we propose the SAM-Guided Diffusion Fusion Network (SGDFuse), a novel framework designed as a principled and effective implementation of this SGG paradigm. However, We find that simply combining these components in a unified, end-to-end framework leads to a significant task conflict—the network struggles to simultaneously optimize low-level, cross-modal feature alignment and high-level, iterative semantic generation. To resolve this, our framework employs a deliberate two-stage design to decouple these conflicting tasks. The first stage is dedicated solely to multi-modal feature extraction and fusion. We design a Multi-Scale Feature Enhancement Module (MSFEM) and a Transformer Block to model the local structures of the infrared image and the global semantics of the visible image, respectively. This stage generates a robust structural prior ($F_{1}$), rather than a final image. The second stage is then oriented towards task-oriented optimization and high-fidelity reconstruction. By receiving $F_{1}$ as its foundation, the learning task for the conditional diffusion model is significantly simplified. It no longer needs to learn how to fuse from scratch but can instead concentrate its powerful generative capabilities on refinement and semantic enhancement. To achieve this, we introduce a holistic semantic guidance system. At the Input, we leverage SAM to generate rich semantic prior masks for both the infrared and visible images. These masks are concatenated with the structural prior ($F_{1}$) to form a five-channel, task-aware input. This explicit channel-wise conditioning is the most direct method to provide dense spatial guidance. During the Process, the deep modeling of this spatio-semantic information is handled by the subsequent denoising U-Net , which is trained to learn the complex correlations between the structural channels and the mask channels. At the Output, this guidance is further enforced by a Hierarchical Feature Aggregation Head (HFAH) and a novel Mask-Guided Loss Function. Through this complete ”Input-Process-Output” integration, our method establishes an effective synergy between structural understanding and detail generation . As our key ablation study (Figure 3) demonstrates, this design significantly boosts the semantic integrity, visual quality, and adaptability for downstream tasks of the resulting fused image.
Figure 3: Visual Ablation for the Necessity of the Semantic-Guided Generation Paradigm. The red boxes show zoomed-in detail regions.
@@
    We propose a new ’Semantic-Guided Generation’ (SGG) paradigm for image fusion. This paradigm conceptually reframes the task from ’pixel reorganization’ to ’semantically-controlled generation’ to fundamentally address the long-standing ’semantic blindness’ problem in the field.
@@
    We propose the SGDFuse network, a novel and principled two-stage framework that serves as the first effective implementation of the SGG paradigm. Its design strategically decouples structural understanding (Stage I) from semantic generation (Stage II), successfully resolving the inherent ’task conflict’ in this new paradigm.
@@
    We design a holistic Input-Process-Output guidance system to operationalize the SGG framework, which includes explicit SAM priors at the input and is critically reinforced by a Mask-Guided Loss function ($L_{stage2}$) at the output, ensuring superior preservation of key semantic information.
@@ 2.2 SAM and Its Applications
The Segment Anything Model (SAM) [13], as a powerful pretrained model, has become a current research hotspot in the field of image processing due to its zero-shot generalization capability in image segmentation and object detection tasks [24]. SAM can provide high-precision semantic masks across various vision tasks and has been widely applied in low-level vision tasks such as dehazing [25], and low-light image enhancement [26], as well as high-level tasks like object detection, demonstrating outstanding performance. By deeply understanding image content, SAM effectively captures fine details and accurately delineates target regions. Therefore, SAM is particularly suitable for infrared and visible image fusion tasks, we leverage SAM-generated segmentation masks as explicit, high-level semantic priors to guide the fusion process.
@@ 3.1 Motivation
Our primary and core motivation is to inject explicit semantic understanding capabilities into the fusion process. We are no longer satisfied with the indiscriminate processing of low-level features by existing methods, but rather seek a novel mechanism that enables the model to truly ”understand” the image and distinguish ”target” from ”background”. For this, we introduce SAM. The high-quality semantic masks it generates can provide the fusion process with unprecedented, strong semantic priors concerning scene structure and object boundaries, which becomes our ideal tool for solving the core problem of ”semantic blindness”. However, high-level semantic guidance alone is insufficient. To accurately translate this precise semantic intent—for instance, ’to preserve the thermal features of a target while enhancing the texture of the background’—into a high-quality, pixel-level output, one must rely on an equally powerful high-fidelity generative framework. Traditional reconstruction methods are prone to generating artifacts and detail loss when handling such complex conditions, which severely undermines the efficacy of semantic guidance. Therefore, our second motivation is to employ a Diffusion Model, renowned for its superior generative capabilities, as the backbone for image reconstruction. This ensures that the high-level semantic commands can be rendered losslessly and with maximum fidelity.
The core of our paradigm is the Deep Coupling of semantic understanding and image generation, rather than treating them as isolated steps. Achieving this, however, required solving two fundamental design challenges, which in turn form our core methodological innovations. First, we must resolve the inherent ”task conflict” between low-level fusion and high-level generation. A naive end-to-end model struggles to simultaneously learn both. Therefore, our first core design choice is a deliberate two-stage architecture. Stage I is dedicated to generating a robust structural prior ($F_{1}$), which significantly simplifies the learning task for Stage II by allowing it to focus only on high-fidelity, semantic-aware refinement. Second, we must move beyond simplistic conditioning to achieve deep semantic guidance. Our second core design choice is a holistic, ”Input-Process-Output” guidance system . We establish an explicit spatial prior at the Input (via concatenation), but this is merely the first step of our complete guidance loop. we use SAM masks as an explicit spatial prior at the Input, which is then deeply modeled by the denoising network (Process), and rigorously enforced at the Output by our novel Mask-Guided Loss ($L_{stage2}$). This complete system, comprising the decoupled architecture and the closed-loop guidance, is the ”Deep Coupling” we seek. It endows the generation process with Precise Semantic Directivity, enabling the model to intelligently preserve thermal targets while reconstructing texture details based on semantic importance. Ultimately, SGDFuse unifies the contradictory objectives of semantic integrity and visual quality, producing fused images that are both visually compelling and highly effective for downstream tasks.
3.2 Network Architecture
In this section, we provide a detailed description of the proposed image fusion framework, SGDFuse. This fusion framework adopts a two-stage architecture, as illustrated in Figure 4, aiming to generate high-quality fused images with structural fidelity and semantic consistency.
@@ 3.3 Multi-Scale Feature Enhancement Module (MSFEM)
As shown in Figure 4(c), the MSFEM employs a multi-branch parallel structure, where each branch utilizes convolutional operations with different receptive fields (e.g., $\times1$, $\times3$, $\times5$, $\times7$) to capture structural features ranging from local to medium scales. Specifically, given an input image, feature representations are first extracted separately by four sets of convolutional kernels:
@@ 3.5 Hierarchical Feature Aggregation Head (HFAH)
To further enhance feature representation in the decoding stage, we introduce a Hierarchical Feature Aggregation Head (HFAH) into the decoder path of the denoising network, as shown in Figure 5. HFAH integrates multi-level decoded features and incorporates a spatial attention mechanism to jointly optimize structural detail and semantic consistency. Specifically, all weighted hierarchical features are concatenated and passed through a fusion head to generate the final fused image $I_{f}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$, which exhibits clearer structural boundaries and improved detail preservation. The fusion head consists of multiple $\times3$ convolutional layers designed to map the high-dimensional, multi-scale diffusion features into a final three-channel fused image. A Tanh activation function is applied at the output to enhance texture continuity and fine detail expression. Benefiting from HFAH’s precise modeling of structural edges and semantic regions, this module significantly improves the boundary fidelity and fine-grained reconstruction quality of the final fused image.
@@ 3.6.2 Stage II: Diffusion Optimization Training
4 EXPERIMENTS
4.1 Experimental Settings
4.1.1 Datasets
[…]
4.1.2 Evaluation Metrics
4.1.3 Training Details
4.2 Generalization Comparison Experiments
To evaluate the generalization capability of our network, we compare it against the following representative image fusion methods: U2Fusion [1], TarDAL [31], PIAFusion [10], IGNet [34], CDDFuse [35], Dif-fusion [5], PFCFuse [36], SHIP [24], EMMA [27], Text-DiFuse [11], DAFusion [15], MaeFuse [6], and SAGE [12]. These SOTA baselines cover a wide range of architectures, including recent diffusion-based models (e.g., Dif-fusion [5], Text-DiFuse [11]) and state-of-the-art methods based on Transformer architectures (e.g., MaeFuse [6] and PFCFuse [36]).
@@ 4.2 Generalization Comparison Experiments
4.2.1 Results and Analysis on the MSRS Dataset
Figure 6 shows qualitative comparisons under diverse conditions, including varying illumination, multiple targets, and day/night scenes. While U2Fusion and PLAFusion exhibit blurred details and brightness inconsistency, IGNet, CDDFuse, and MaeFuse partially enhance thermal targets but often lack texture fidelity or introduce artifacts. In contrast, our method effectively preserves thermal saliency, enhances structural clarity, and maintains global luminance balance, achieving superior perceptual quality and semantic consistency.
4.2.2 Results and Analysis on the M3FD Dataset
Figure 7 shows qualitative comparisons on challenging scenes from the M3FD dataset, including nighttime roads, uneven lighting, low-contrast IR targets, and complex textures. Methods like U2Fusion and TarDAL suffer from blurred structures and missing textures, while IGNet, PLAFusion, and CDDFuse offer partial improvements but still lack edge continuity and luminance balance. MaeFuse and SAGE improve global coherence but exhibit limitations in detail preservation. In contrast, our method consistently restores thermal contours and visible textures with natural brightness and clear contrast. Especially in scenes with weak IR signals and cluttered backgrounds, SGDFuse achieves accurate foreground–background separation, highlighting its superiority in structural refinement and semantic consistency.